OpenStudy (anonymous):

I have a quiz on the chain rule today and I don't understand anything about it. Can someone please help?

OpenStudy (anonymous):

chain rule : dy/dx = (dy/du) x (du/dx)

OpenStudy (anonymous):

eg. y=(x^2 +1)^6 Then dy/dx = (dy/du) x (du/dx) = 6(x^2 +1 )^5 X 2x = 12x(x^2+1)^5 hopes this helps! =D

OpenStudy (anonymous):

Thanks a lot!!

OpenStudy (anonymous):

no worries!